OpenStudy (anonymous):

Just out of interest - i found a surprising result in complex number theory - evaluate i^i using the identity cos x + i sin x = e^ix where x is in radians.

OpenStudy (anonymous):

you let x = pi/2 so e ^(i pi) = cos (pi/2) + i sin (pi/2)

OpenStudy (anonymous):

that last line should be e ^(i pi/2) = cos (pi/2) + i sin (pi/2)

OpenStudy (anonymous):

so e ^(i pi) = 0 + i e ^(i pi/2)^1 = i^i i^i = e^(-pi/2) = approx 0.208 - a real number!

OpenStudy (anonymous):

yes you get \[e^{-\frac{\pi}{2}}\]

OpenStudy (anonymous):

but of course that is the "principle branch" because there are infinitely many

OpenStudy (anonymous):

since the representation is not unique i.e. you can use \[i=e^{\frac{5\pi}{2}}\]as well

OpenStudy (anonymous):

or more generally \[i=e^{\frac{(2n-1)\pi}{2}}\]

OpenStudy (anonymous):

right : pi/2 + 2npi

OpenStudy (anonymous):

but it is cool that all are real!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and of course i made a typo both times by omitting the i!

OpenStudy (anonymous):

oh well!!

OpenStudy (anonymous):

also... you can use \[i^i=e^{i\ln(i)}\] and by some miracle it turns out that \[\ln(i)=\frac{\pi i}{2}\] giving the same result

OpenStudy (anonymous):

well not really a miracle it says the same thing. \[log(z)=\log(r)+i\log(\theta)\]

OpenStudy (anonymous):

great stuff

OpenStudy (anonymous):

wish i remembered more of it that is for sure. mostly gone from my brain. someone actually asked this as a question here not too long ago

OpenStudy (anonymous):

right - math never fails to surprise. and much is still unknown - more than is known, i suspect.

OpenStudy (anonymous):

i had a professor once who believed that just like rational numbers have measure 0 in the reals so decidable problems have measure 0

OpenStudy (anonymous):

interesting - well - back to the mundane - my wife wants to go shopping!! see u

OpenStudy (anonymous):

yeah time for market for me as well.

OpenStudy (anonymous):

i'm sorry i missed this, very interesting!